OpenStudy (jennyyy666):

seven is at most the quotient of a number d and -5

jhonyy9 (jhonyy9):

do you know what mean a quotient of a number d ?

jhonyy9 (jhonyy9):

for example than you divide a divide b = c so what is the quotient ?

jhonyy9 (jhonyy9):

@Will.H

OpenStudy (will.h):

she's not responding

OpenStudy (will.h):

in jhonny's case the quotient is c

jhonyy9 (jhonyy9):

yes probably like getting directly the right answer what is opposite by the rules of OS.

jhonyy9 (jhonyy9):

jenny please collaborate

OpenStudy (jennyyy666):

I don't know what the quotient means. I am good at putting equations together but this problem is just confusing for me.

OpenStudy (jennyyy666):

I put that for my answer and its saying it is wrong.

OpenStudy (phi):

Seven is at most x means 7 is no more than x or 7 is less than or equal to x or \[ 7 \le x \] in this case, x is the quotient of d and -5 i.e. d/-5 (this can be confusing, and I would not say it like that) anyway the answer is \[ 7 \le \frac{d}{-5} \]

OpenStudy (mathstudent55):

The quotient is the result of a division. The quotient of 20 and 2 is the result of the division 20/2 which is 10. In your case, the quotient of a number d and -5 is the result of the division of d by -5, which is \(\dfrac{d}{-5} \)